# LiveMCPBench: Can Agents Navigate an Ocean of MCP Tools?

### v1

@@
Abstract
With the rapid development of Model Context Protocol (MCP), the number of MCP servers has surpassed 10,000. However, existing MCP benchmarks are limited to single-server settings with only a few tools, hindering effective evaluation of agent capabilities in large-scale, real-world scenarios. To address this limitation, we present LiveMCPBench, the first comprehensive benchmark comprising 95 real-world tasks grounded in the MCP ecosystem, designed to evaluate LLM agents at scale across diverse servers. To support a scalable and reproducible evaluation pipeline in large-scale MCP environments, we curate LiveMCPTool, a diverse and readily deployable collection of 70 MCP servers and 527 tools. Furthermore, we introduce LiveMCPEval, an LLM-as-a-Judge framework that enables automated and adaptive evaluation in dynamic, time-varying task environments, achieving 81% agreement with human reviewers. Finally, we propose the MCP Copilot Agent, a multi-step agent that routes tools for dynamic planning and executes tools for API interaction across the entire LiveMCPTool suite. Our evaluation covers 10 leading models, with the best-performing model (Claude-Sonnet-4) reaching a 78.95% success rate. However, we observe large performance variance across models, and several widely-used models perform poorly in LiveMCPBench’s complex, tool-rich environments. Overall, LiveMCPBench offers the first unified framework for benchmarking LLM agents in realistic, tool-rich, and dynamic MCP environments, laying a solid foundation for scalable and reproducible research on agent capabilities. Our code and data will be publicly available at https://icip-cas.github.io/LiveMCPBench.
Introduction
@@
(b) MCP Benchmarks: Small-scale MCP servers.
(c) LiveMCPBench: Real & Large-scale MCP servers.
Figure 1: Comparison between LiveMCPBench and existing Tool-Use and MCP benchmarks: We focus on real-world task resolution using large-scale MCP toolset.
Figure 2: LiveMCPBench comprises four parts: (1) Diverse Daily Task, a collection of everyday tasks; (2) LiveMCPTool, a large-scale, out-of-the-box MCP toolset; (3) MCP Copilot Agent, a ReACT-based agent system capable of retrieval and multi-turn tool invocation across the MCP toolset; (4) LiveMCPEval, an automated LLM-as-a-Judge evaluation system for accurately assessing online, time-varying tasks.
Recent years have witnessed remarkable progress in tool-use agents powered by large language models (LLMs), demonstrating promising potential as a pathway towards artificial general intelligence (Qu et al. 2025; Wang et al. 2024). As model capabilities advance and application scenarios expand, enabling LLMs to effectively invoke external tools has emerged as a critical research direction for enhancing their generalization and real-world task execution abilities. Notably, with the widespread adoption of the Model Context Protocol (Anthropic 2024, MCP), an increasing number of real-world tools now expose their functionalities through standardized contextual interfaces, forming a vast ecosystem encompassing more than 10,000 MCP servers (Hou et al. 2025). Concurrently, certain pretrained models have begun to directly learn interaction patterns with MCP (Qwen 2025), further accelerating the evolution of tool-use agents.
However, existing tool-use benchmarks predominantly rely on simulate API interfaces (see Figure 1), which suffer from a fundamental limitation. API interfaces exhibit high instability—For instance, 55.6% of APIs in ToolBench (Qin et al. 2024) have become unavailable (Guo et al. 2024), forcing evaluation frameworks (e.g., API-Bank (Li et al. 2023)) to resort to simplified simulated tools, significantly compromising task authenticity and challenge. And the emergence of MCP provides a stable tool call interface. For the few MCP benchmarks (e.g., MCPBench (Luo et al. 2025) and MCPEval (Liu et al. 2025)), their experimental scales remain severely limited, typically involving only a small number of MCP servers (about 10), failing to reflect agents’ generalization and decision-making capabilities in a large-scale toolset.
Consequently, two pivotal questions remain underexplored: (1) How can optimal planning and retrieval be achieved in large-scale MCP toolset to accomplish real-world tasks? (2) Can an LLM-driven agent, trained for tool invocation, exhibit meta-tool-learning capabilities—i.e., autonomously explore and compose tools from real-world toolset to complete tasks?
To address these questions, we present LiveMCPBench, a multi-domain, multi-step MCP tool-use benchmark. LiveMCPBench focuses on practical, everyday tasks across six domains, comprising 95 high-quality tasks solvable via MCP tools. To lower the deployment barrier and avoid complex multi-API-key dependencies, we introduce LiveMCPTool, a ready-to-use, fully functional MCP toolset encompassing 70 MCP servers and 527 tools, significantly improving reproducibility and reducing research overhead.
Furthermore, evaluating agent systems in real-world MCP environments presents two unique challenges: (1) Task dynamism (e.g., time-sensitive answers for news summarization) and (2) Solution diversity, multiple tool combinations can solve the same task. Consequently, traditional metrics based on tool-matching accuracy in API-Bank become inadequate. We thus design LiveMCPEval, an LLM-as-a-Judge (Zheng et al. 2023) evaluation system capable of automatically assessing multi-turn tool invocation trajectories.
Finally, the server-tool architecture of MCP fundamentally differs from traditional parallel API structures, rendering existing tool-use agent methods inapplicable in MCP environments. To evaluate planning and retrieval strategies in large-scale MCP toolset and verify whether LLMs possess meta-tool-learning capabilities, we propose MCP Copilot Agent—an agent integrating reasoning and acting (Yao et al. 2023, ReACT) strategy with tool retrieval and multi-step invocation abilities, capable of dynamically responding to environmental changes. Our evaluation of 10 frontier models reveals that the Claude-Sonnet-4 agent achieves a 78.95% task success rate, demonstrating strong meta-tool-learning capabilities. We further conduct human annotation on the agent trajectory and measure the human agreement rate across different evaluator models, with DeepSeek-V3 achieving an agreement rate of 81.05%, validating the reliability of our evaluation method.
@@
  • We propose LiveMCPBench, the first evaluation framework leveraging large-scale MCP toolset for everyday tasks.
  • We evaluate the frontier models using LiveMCPBench, revealing their limitations in meta-tool-learning.
  • We analyze the agent trajectories, identifying key bottlenecks in multi-tool collaboration and providing insights for future improvements.
LiveMCPBench
@@
Type Stable Servers Tools Plug & Play Type Time-Varying Evaluation
@@
MCPBench (Luo et al. 2025) MCP 10 10 Real Rule
MCP-RADAR (Gao et al. 2025) MCP 9 42 Real Rule
MCP-Zero (Fei, Zheng, and Feng 2025) MCP 308 2,797 - - -
MCPEval (Liu et al. 2025) MCP 12 77 Syn. LLM
LiveMCPBench (ours) MCP 70 527 Real LLM
Tool Benchmarks
ToolAlpaca (Tang et al. 2023) Simulate 50 426 Syn. Rule
ToolBench (Qin et al. 2024) API 49 16,464 Syn. LLM
API-Bank (Li et al. 2023) Simulate 1,000 2,138 Real Rule
Table 1: Comparison with existing MCP and Tool Benchmarks. Plug & Play indicates whether the toolset can be used directly without additional API keys. Time-Varying denotes whether the task output changes over time. Note that MCP-Zero is not a benchmark, but contains a toolset. And, it is not directly usable, as many of its tools require API keys. Syn. means synthetic.
We present LiveMCPBench (see Figure 2), a novel benchmark designed to evaluate the capability of agent systems in retrieving appropriate tools from a large-scale MCP toolset to accomplish general-purpose everyday tasks. The construction of such a benchmark necessitates addressing three fundamental challenges:
  • How to construct representative daily tasks requiring multi-step tool use?
  • How to collect a large, redundant yet functionally complete MCP toolset?
  • How to automatically evaluate performance on evolving online tasks?
In this section, we first elaborate on our task construction process. Then, we outline the collection process of the LiveMCPTool. Next, we introduce LiveMCPEval, an LLM-as-a-Judge evaluation framework designed for robust and scalable assessment. Finally, we describe MCP Copilot Agent as our baseline approach.
Task Construction
To advance the development of practical agents for real-world applications, we create a diverse set of tasks grounded in everyday scenarios, spanning six key domains: Office (e.g., spreadsheet analysis), Lifestyle (e.g., news retrieval), Leisure (e.g., video game inquiries), Finance (e.g., stock price monitoring), Travel (e.g., ticket search), and Shopping (e.g., product recommendations). These scenarios were carefully selected to embody three critical characteristics. (1) Time-varying: Tasks exhibit time-sensitive outcomes; (2) Long-horizon: Tasks require multiple tools to complete; (3) Genuine utility: Tasks address authentic user needs.
The annotation process employed a rigorous two-stage methodology involving two groups of computer science students serving as task proposers and validators. Proposers first generated scenario-specific tasks based on personal experience, with LLM-assisted ideation permitted but strictly vetted for authenticity. Each proposer then interacted with our toolset to complete their proposed task, meticulously annotating key points to preserve the task’s compositional depth. Validators subsequently scrutinized both the task design and corresponding toolchain invocations, eliminating duplicates while enforcing quality standards. This iterative pipeline yielded 95 high-fidelity daily tasks (see Appendix C for annotation principles and distribution statistics).
Model Office Leisure Travel Lifestyle Finance Shopping Overall (%)
Claude-Sonnet-4-20250514 90.32 64.29 75.00 80.00 78.57 66.67 78.95
Claude-Opus-4-20250514 80.65 64.29 66.67 86.67 64.29 33.33 70.53
DeepSeek-R1-0528 41.94 50.00 58.33 46.67 50.00 55.56 48.42
Qwen3-235B-A22B 54.84 35.71 41.67 53.33 50.00 44.44 48.42
GPT-4.1-Mini 45.16 50.00 50.00 46.67 42.86 22.22 44.21
Qwen2.5-72B-Instruct 35.48 35.71 50.00 40.00 57.14 55.56 43.16
DeepSeek-V3-0324 41.94 42.86 50.00 40.00 28.57 55.56 42.11
Gemini-2.5-Pro 48.39 28.57 16.67 60.00 57.14 11.11 41.05
GPT-4.1 51.61 28.57 25.00 46.67 35.71 22.22 38.95
Qwen3-32B 29.03 14.29 25.00 53.33 28.57 33.33 30.53
Table 2: Task success rate results for the frontier models. Evaluation using Deepseek-V3.
LiveMCPTool Collection
@@
While prior study (Hou et al. 2025) suggests the existence of over 10,000 MCP servers, curating a practical and accessible toolset remains nontrivial due to critical usability constraints. The predominant challenge stems from dependency fragmentation: the majority of MCP servers necessitate proprietary API keys or integrations with third-party services, rendering them impractical for a standardized toolset. To address this, we introduce a rigorously validated methodology for constructing a high-quality, dependency-free MCP toolset—prioritizing reproducibility and broad applicability. Our approach first aggregates 5,588 server configurations from mcp.so, then systematically filters out key-dependent servers to eliminate access barriers.
@@
Beyond accessibility, we ensure the toolset’s representativeness through structured curation and expert annotation. Tools are taxonomically organized into five functional categories (Discovery, Visualization, File Access, Location, and Miscellaneous), followed by manual vetting to exclude low-quality implementations. This two-stage pipeline yields 70 MCP servers providing 527 tools, each verified for standalone functionality and categorical relevance. By decoupling the toolset from external dependencies, our collection establishes the first reproducible toolset for large-scale MCP performance analysis (see Appendix D for distribution details).
LiveMCPEval
Automated evaluation of trajectories generated by the agent is essential for benchmarking task performance. However, achieving robust automated evaluation remains challenging due to several factors: (1) Time-varying nature of daily tasks, (2) Inherent instability of MCP tool outputs caused by its online dynamics, and (3) Diversity of trajectories resulting from different tool combinations that can accomplish the same task.
To address these challenges, we employ an LLM-as-a-Judge system, leveraging the adaptability of LLMs to dynamically assess task completion based on tool usage patterns and feedback. While dynamic tasks may exhibit variability, they often share a set of critical subtasks or key points that must be fulfilled. Incorporating these key points—whether manually annotated or automatically extracted by LLM—improves the accuracy of the LLM-as-a-Judge system. In our framework, all tasks are annotated with a verified set of key points to ensure a reliable evaluation. Specifically, given a task $T$, a set of key points $P$, agent’s execution trajectory $A$ with retrieval and tool-call sequences and descriptions of all tools used $D$, the evaluator performs binary classification to determine the outcome $\mathcal{O}$ as either “Success” or “Failure”:
$\mathcal{O}=Evaluator(T,P,A,D)$ (1)
Therefore, we use the success rate as the primary metric. Finally, we systematically compare LiveMCPBench with existing benchmarks in Table 1, highlighting their fundamental differences.
MCP Copilot Agent
Due to the dynamic nature of daily tasks and the inherent uncertainty in retrieval systems, a fixed pipeline cannot be effectively employed for tool retrieval and invocation in LiveMCPTool. Instead, we require agents to dynamically adapt to environmental changes. To model this dynamic tool retrieval and invocation process, we formulate it as a Partially Observable Markov Decision Process (Silver and Veness 2010, POMDP), as the agent can only make decisions based on the textual descriptions of retrieved tools and feedback from tool execution.
We characterize the toolset environment using the following components: (1) Hidden state space $\mathcal{S}$; (2) Observation space $\mathcal{O}$ containing the descriptions of retrieved tools and feedback from tools; (3) Language action space $\mathcal{A}$, including three key actions—route, execute, and response—along with their associated descriptions; (4) State transition $\mathcal{T}:\mathcal{S}_{t}\times\mathcal{A}\rightarrow\mathcal{S}_{t+1}$; (5) Terminal reward $\mathcal{R}:\mathcal{S}\rightarrow\mathbb{R}$ quantifying task completion.
Our agent implementation is based on the ReACT framework. For the route tool, we adopt a retrieval strategy inspired by MCP-Zero (Fei, Zheng, and Feng 2025), where tool prioritization is determined by a weighted combination of server description similarity and tool description similarity.
Experiments and Results
Setup
We evaluate 10 frontier models: Claude-Opus-4 and Claude-Sonnet-4 (Anthropic 2025), GPT-4.1 and GPT-4.1-Mini (Openai 2025), Gemini-2.5-Pro (Google 2025), Deepseek-V3 and Deepseek-R1 (DeepSeek-AI 2025), Qwen3-235B-A22B and Qwen3-32B (Qwen 2025), and Qwen2.5-72B-Instruct (Qwen et al. 2025). For assessment, we employ Deepseek-V3 as our primary evaluation model. Detailed implementation and computational resources are documented in Appendix E to ensure reproducibility.
Main Results
We show the task success rates for different models in Table 2. We can see that:
  1. 1.
    Meta-Tool-Learning Capabilities in Claude Models. The Claude series demonstrates remarkable meta-tool-learning proficiency, with Claude-Sonnet-4 and Claude-Opus-4 achieving success rates of 78.95% and 70.53% respectively. These results indicate their superior ability to effectively explore and combine tools from a large-scale toolset to accomplish complex real-world tasks.
  2. 2.
    Performance Variance Among Models. We observe substantial performance gap across frontier models. While most contemporary models achieve only 30%–50% task success rates, the Claude series shows significantly superior performance. This performance gap suggests fundamental limitations in the meta-tool-learning capabilities of other models.
  3. 3.
    Domain-Specific Superiority of Claude Models. The Claude series exhibits particularly dominant performance in Office and Lifestyle scenarios, outperforming other models by more than 30%. This substantial advantage highlights Claude models’ unique strengths and adaptability in these specific domains.
Figure 3: Correlation between model evaluation performance and human agreement rates across multiple evaluators: The analysis of annotated agent trajectories from Claude-Opus-4 and Claude-Sonnet-4 (color indicates evaluator variant).
To validate the reliability of LiveMCPEval’s automatic evaluation, we conducted human annotation of the execution trajectories for the top-performing models (Claude-Sonnet-4 and Claude-Opus-4). We systematically tested all models used in the baselines and calculated human agreement rates, with results presented in Figure 3.
  1. 1.
    LiveMCPEval demonstrates high accuracy under appropriate model conditions. Our experiments show that Deepseek-V3 achieves an average human agreement rate of 78.95%, validating the reliability of our evaluation framework. Additionally, GPT-4.1 Mini and Qwen2.5-72B-Instruct exhibit comparable performance, with human agreement rates around 75%, making them viable alternative models for accurate assessment.
  2. 2.
    Certain models prove less suitable for evaluation tasks. Notably, advanced reasoning models such as Deepseek-R1, Claude-Opus-4, and Qwen3-32B exhibit lower human agreement rates (60%-70%). We hypothesize that this limitation stems from their reduced ability to process long trajectory inputs.
Figure 4: Comparison of human agreement rates across different models when evaluated using human-annotated versus LLM-generated key points.
To investigate the generalizability of LiveMCPEval, we conducted experiments on Claude-Sonnet-4’s trajectories by evaluating how LLM-generated key points affect human agreement rates, with results visualized in Figure 4 (Detaied examples in Appendix F). Our key findings reveal:
  1. 1.
    Generalizability of LiveMCPEval. The framework demonstrates strong generalizability across models. Notably, even without human-annotated references, the majority of evaluated models achieved improved human agreement rates through automatically generated key points.
  2. 2.
    Optimal Utilization of Human-Annotated Key Points. Deepseek-V3 exhibits superior capability in leveraging human-annotated key points compared to other models. This observation provides important implications for model selection in scenarios where human-annotated references are available.
Analysis
Model Steps Tools excute route Overall (%)
Claude-Sonnet-4 20.09 2.71 5.59 2.98 78.95
Claude-Opus-4 25.53 3.40 6.93 4.35 70.53
Qwen3-235B 16.76 1.59 5.12 1.77 48.42
DeepSeek-R1 10.33 1.24 2.11 2.00 48.42
GPT-4.1-Mini 10.89 1.37 2.71 1.65 44.21
Qwen2.5-72B 11.22 1.31 2.80 1.38 43.16
DeepSeek-V3 8.33 1.01 1.29 1.41 42.11
Gemini-2.5-Pro 8.08 0.99 1.46 1.35 41.05
GPT-4.1 9.03 1.31 1.72 1.64 38.95
Qwen3-32B 9.99 1.16 2.31 1.19 30.53

### v2

@@
Abstract.
Model Context Protocol (MCP) has become a key infrastructure for connecting LLMs with external tools, scaling to 10,000+ MCP servers with diverse tools. Unfortunately, there is still a large gap between real-world MCP usage and current evaluation: they typically assume single-server settings and directly inject tools into the model’s context, bypassing the challenges of large-scale retrieval and multi-tool composition. To bridge this gap, we propose LiveMCPBench, which evaluates 95 real-world daily tasks explicitly constructed to stress diverse tools and scaled multi-server routing. The benchmark includes a ready-to-deploy tool suite of 70 servers with 527 tools, ensuring reproducibility without scattered API configuration. We further introduce an LLM-as-a-Judge evaluation framework that directly verifies task outcomes, handling dynamic data sources and multiple valid solution paths. We benchmark 12 state-of-the-art LLMs and observe a substantial performance gap: while Claude-Sonnet-4 reaches 78.95% task success, most models achieve only 30–50%. Our analysis reveals that active tool composition strongly correlates with task success, whereas retrieval errors account for nearly half of all failures—highlighting retrieval as the dominant bottleneck. Together, these results provide the first large-scale, reproducible diagnosis of MCP agent capabilities and point towards future research on improving retrieval robustness and encouraging effective tool composition. Our code and data are publicly available at https://icip-cas.github.io/LiveMCPBench.
Model Context Protocol, Scalable MCP-use, Benchmark
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologies Artificial intelligenceccs: General and reference Evaluation
1. Introduction
Tool-use agents powered by large language models (LLMs) are emerging as a critical step toward general intelligence (Qu et al., 2025; Wang et al., 2024). The Model Context Protocol (Anthropic, 2024, MCP) has rapidly become the core infrastructure for connecting LLMs with external tools (Ehtesham et al., 2025). Its ecosystem is rapidly scaling to more than 10,000 servers with diverse functionalities (Hou et al., 2025). Meanwhile, pretrained models are being equipped with the ability to invoke MCP servers directly (Qwen, 2025). MCP has greatly expanded the boundaries of tool-use agents, enabling them to tackle increasingly complex real-world tasks (Ray, 2025). As a result, the ability to use MCP effectively has become a key criterion of an agent’s capability.
Despite the rapid expansion of MCP, existing benchmarks fall short of capturing its real-world challenges, as shown in Figure 1. Traditional tool-use benchmarks such as API-Bank (Li et al., 2023) and ToolBench (Qin et al., 2024) rely on simulated or unstable APIs, resulting in less realistic and unreliable tasks. For example, 55.6% of APIs in ToolBench are no longer available (Guo et al., 2024). In contrast, MCP provides a more stable invocation interface through a package-like management model, where servers can be reliably maintained and provide both local and online functionality. MCP adopts a server-tool architecture, where each server exposes multiple related tools with contextual descriptions, providing richer semantics. This structure requires agents to develop retrieval and composition abilities that go beyond conventional tool search. However, current MCP benchmarks remain narrow in scope. For example, MCPBench (Luo et al., 2025) covers only 10 servers and directly injects tools into the LLM’s context, bypassing large-scale retrieval and multi-tool composition entirely. Furthermore, most existing MCP benchmarks are still conducted in static environments, making them unable to evaluate agents’ performance in dynamic, large-scale MCP ecosystems.
These limitations lead to two central research questions:
  • RQ1: How can we systematically evaluate an agent’s retrieval and composition capabilities in large-scale, diverse MCP ecosystems?
  • RQ2: How can we design evaluation methodologies that are both scalable and reproducible under dynamic, real-world settings?
To address these questions, we introduce LiveMCPBench, a benchmark for realistic, large-scale MCP evaluation. LiveMCPBench covers six practical domains and consists of 95 multi-step, daily tasks that explicitly test retrieval and composition abilities. We focus on daily scenarios because most MCP servers are designed to provide up-to-date information (e.g., news), making them a natural and challenging frontier for assessing MCP-use capabilities. To facilitate reproducibility and ease of use, we provide LiveMCPTool, a curated collection of 70 MCP servers with 527 tools, packaged as a ready-to-use environment that removes the need for configuring numerous scattered API keys. For evaluation, we propose LiveMCPEval, an automatic evaluation framework based on the LLM-as-a-Judge (Zheng et al., 2023), enabling scalable assessment even under dynamic data sources (e.g., evolving weather or news) and supporting multiple valid solution trajectories. Crucially, it supports scalable and open-ended evaluation, allowing seamless expansion as new MCP servers and tasks are introduced, without costly manual annotation.
@@ 1. Introduction
(b) MCP Benchmarks: Real but Small-scale MCPs.
(c) LiveMCPBench: Real & Large-scale MCPs.
Figure 1. LiveMCPBench focuses on daily task resolution using a large-scale MCP toolset.
Building on LiveMCPBench, we develop MCP Copilot Agent, a ReACT agent (Yao et al., 2023) capable of retrieving and composing tools across a large MCP toolset. We evaluate 12 leading LLMs on the benchmark and find a substantial performance gap: while Claude-Sonnet-4 achieves a 78.95% task success rate, the majority of widely used frontier models remain in the 30–-50% range. By analyzing agent behaviors, we observe that active tool composition strongly correlates with task success, indicating that models capable of flexibly combining tools are more reliable. A detailed error analysis further shows that nearly half of all failures stem from tool retrieval errors, establishing retrieval as the primary bottleneck in large-scale MCP usage. These findings highlight the urgent need for future systems that jointly improve tool composition and mitigate retrieval errors. To further validate our evaluation pipeline, we conduct human annotation of agent trajectories and assess human–model agreement. DeepSeek-V3 attains an agreement rate of 81.05%, demonstrating the reliability of our evaluation protocol.
@@
  • We introduce LiveMCPBench, a benchmark with 95 multi-step daily tasks that test agents’ large-scale retrieval and multi-tool composition abilities in large-scale MCP ecosystems.
  • We provide a reproducible toolset (70 MCP servers, 527 tools) together with an LLM-as-a-Judge automatic evaluation framework, enabling scalable, automated, and human-aligned assessment in dynamic environments.
  • Our empirical study offers a large-scale diagnosis of MCP agent capabilities, highlighting tool retrieval and composition as unsolved challenges and laying the groundwork for future research.
2. LiveMCPBench
Figure 2. LiveMCPBench comprises four parts: (1) Diverse Daily Task, a collection of 95 daily tasks; (2) LiveMCPTool, a large-scale MCP toolset; (3) MCP Copilot Agent, a ReACT-based agent supporting retrieval and multi-turn tool invocation across the MCP toolset; (4) LiveMCPEval, an automated LLM-as-a-Judge evaluation system for accurately assessing dynamic tasks.
We present LiveMCPBench (Figure 2), a benchmark designed to evaluate the ability of agent systems to retrieve appropriate tools from a large-scale MCP toolset for accomplishing general-purpose daily tasks. Building such a benchmark requires addressing three key challenges:
  • Designing representative tasks that faithfully capture and expose the unique characteristics and inherent complexities of large-scale MCP ecosystems.
  • Constructing a comprehensive, reproducible toolset that remains both large and functionally complete while being easy to deploy for consistent and fair benchmarking.
  • Automating robust evaluation of agent performance on dynamic, evolving online tasks, where task context and available tools may change over time.
In this section, we first detail our task construction methodology. Then, we describe the collection process of the LiveMCPTool. Next, we introduce LiveMCPEval, an LLM-as-a-Judge evaluation framework that enables robust and scalable assessment. Finally, we build MCP Copilot Agent, which serves as our baseline approach.
2.1. Task Construction
To ensure that LiveMCPBench reflects the practical strengths of MCP servers, we design tasks grounded in everyday scenarios where these servers are most impactful. Since many MCP servers are oriented toward real-time information retrieval (e.g., news, stock prices, weather updates) or local office productivity (e.g., spreadsheet analysis, scheduling), daily-life settings naturally highlight their distinctive capabilities. By anchoring tasks in such contexts, we capture both the immediacy and utility that characterize MCP-driven interactions.
The resulting benchmark spans six daily domains: Office (e.g., spreadsheet analysis), Lifestyle (e.g., news retrieval), Leisure (e.g., video game inquiries), Finance (e.g., stock price monitoring), Travel (e.g., ticket search), and Shopping (e.g., product recommendations). Tasks in these domains are inherently dynamic, often depending on evolving contexts such as monitoring fluctuating markets. They are also long-horizon, as solving them typically requires coordinating multiple tools rather than a single call. At the same time, the tasks retain genuine utility, reflecting realistic and meaningful user needs that go beyond artificial benchmarks.
The task creation process employed a rigorous two-stage methodology involving two groups of computer science students serving as task proposers and validators. Each group consisted of three members to ensure both diversity of perspectives and reliability of annotation. Proposers first generated scenario-specific tasks based on personal experience, with LLM-assisted ideation permitted but strictly vetted for authenticity. Each proposer then interacted with our toolset to complete their proposed task, meticulously annotating key points to preserve the task’s compositional depth. Validators subsequently scrutinized both the task design and corresponding toolchain invocations, eliminating duplicates while enforcing quality standards. In total, proposers initially produced 300 candidate tasks across six domains. After iterative validation and refinement, this pipeline yielded 95 high-fidelity daily tasks (see Appendix E for annotation principles and distribution statistics).
2.2. LiveMCPTool Collection
Table 1. Comparison with existing work. Plug & Play denotes the toolset requires no extra API keys. Dynamic denotes the task depending on evolving contexts. Syn. denotes tasks generated by LLMs. MCP-Zero is not a benchmark, but contains a toolset.
@@ 2.2. LiveMCPTool Collection
Stable Servers Tools Plug & Play Type Number Dynamic Evaluation
Tool-Use Benchmarks
API-Bank (Li et al., 2023) 8 73 Real 314 Rule
ToolBench (Qin et al., 2024) 49 3,451 Syn. 126,486 LLM
$\tau$-bench (Yao et al., 2025) 2 28 Real 165 Rule
@@
MCPBench (Luo et al., 2025) 10 10 Real 911 Rule
MCP-RADAR (Gao et al., 2025) 9 42 Real 300 Rule
MCP-Zero (Fei et al., 2025) 308 2,797 - - - -
MCPEval (Liu et al., 2025) 12 77 Syn. 676 LLM
LiveMCPBench (ours) 70 527 Real 95 LLM
@@
While prior study (Hou et al., 2025) suggests the existence of over 10,000 MCP servers, creating a practical and accessible toolset remains nontrivial due to critical usability constraints. The predominant challenge stems from dependency fragmentation: the majority of MCP servers necessitate proprietary API keys or integrations with third-party services, rendering them impractical for a standardized toolset. To address this, we introduce a rigorously validated methodology for constructing a high-quality, dependency-free MCP toolset—prioritizing reproducibility and broad applicability. Our approach first aggregates 5,588 server configurations from a MCP Marketplace111mcp.so, then systematically filters out key-dependent servers to eliminate access barriers.
@@
Beyond accessibility, we ensure the toolset’s representativeness through structured curation and expert annotation. Tools are taxonomically organized into five functional categories (Discovery, Visualization, File Access, Location, and Miscellaneous), followed by manual vetting to exclude low-quality implementations. To ensure reproducibility, we package all collected tools in a docker environment, providing a stable and consistent tool setup. This two-stage pipeline yields 70 MCP servers providing 527 tools, each verified for standalone functionality and categorical relevance. By decoupling the toolset from external dependencies, our collection establishes a reproducible toolset for large-scale MCP performance analysis (see Appendix F for distribution details).
2.3. LiveMCPEval
Evaluating agent trajectories in dynamic environments is non-trivial, as task outcomes often depend on evolving contexts, and multiple tool invocation trajectories may lead to equally valid solutions. Such properties make traditional metrics, such as exact tool-matching accuracy, inadequate. To address this, we introduce LiveMCPEval, an automatic evaluation framework that leverages an LLM-as-a-Judge system. Instead of relying on fixed ground-truth trajectories, LiveMCPEval directly verifies task completion by considering tool usage patterns and contextual feedback. A central component of this design is the use of key points—critical subtasks or intermediate conditions that must be satisfied for success—which can be either manually annotated or automatically extracted by LLMs (see Appendix H for examples). In our benchmark, all tasks are equipped with a verified set of key points to ensure reliable evaluation. Formally, given a task $T$, a set of key points $P$, an execution trajectory $A$ (including retrievals and tool calls), and tool descriptions $D$, the evaluator performs binary classification to determine task success:
(1) $\mathcal{J}=Evaluator(T,P,A,D)$
where $\mathcal{J}\in\{\text{Success},\text{Failure}\}$. The overall benchmark metric is the success rate across tasks. Beyond accuracy, the framework is designed for openness and scalability: as new MCP servers and tasks emerge, LiveMCPEval enables seamless benchmark expansion without costly manual annotations, ensuring its long-term relevance in MCP ecosystems. A systematic comparison with existing benchmarks is presented in Table 1.
2.4. MCP Copilot Agent
Daily tasks are inherently dynamic, making fixed pipelines for tool use impractical. For instance, always retrieving before executing may fail when outputs must be combined across tools or when re-routing is required after errors. Agents must instead adapt to evolving contexts.
We formulate tool retrieval and invocation as a Partially Observable Markov Decision Process (Silver and Veness, 2010, POMDP), where decisions rely only on tool descriptions and execution feedback. The environment is defined by: (1) hidden states $\mathcal{S}$ for the underlying task; (2) observations $\mathcal{O}$ from retrieved descriptions and feedback; (3) a language action space $\mathcal{A}$; (4) transition $\mathcal{T}:\mathcal{S}_{t}\times\mathcal{A}\rightarrow\mathcal{S}_{t+1}$; (5) reward $\mathcal{R}:\mathcal{S}\rightarrow\mathbb{R}$ for task success. The action space $\mathcal{A}$ consists of three operations: Route, which retrieves $k$ candidate tools (fixed at $k=5$ in our experiments) from the full toolset; Execute, which invokes a selected tool with specified parameters and returns its output as feedback; Response, which terminates the process and delivers the final result to the user. Tool score in Route follows MCP-Zero (Fei et al., 2025):
(2) $\text{score}=(s_{\text{server}}\times s_{\text{tool}})\times\max(s_{\text{server}},s_{\text{tool}})$
where $s_{\text{server}}$ and $s_{\text{tool}}$ denote cosine similarity to server-level and tool-level descriptions. This form emphasizes joint alignment while letting server priors dominate when tool matches are weak.
Our agent builds on ReACT, integrating reasoning and action in a unified loop. Unlike conventional APIs treating tools as isolated black boxes, MCP introduces a server–tool hierarchy that demands structural reasoning and coordination. The MCP Copilot Agent thus adapts flexibly under uncertainty while exploiting MCP’s hierarchical richness.
3. Experiments and Results
3.1. Setup